Determine the Green function and express the solution as a definite integral

-(y''+y)=f(x), y'(0)=0, y(1)=0


1
Expert's answer
2021-08-19T13:04:22-0400

We remind that the Green function is a function "G(x,s)" such that "y(x)=\int_{x_0}^xG(x,s)f(s)ds".

We rewrite the equation as "z=y'" , "-z'-y=f(x)". It can be rewritten as: "\left\{\right.\begin{matrix}
 y'=z \\
 z'=-y-f(x)
\end{matrix}"

It can be rewritten as: "\left(\begin{matrix}
1&0
 \end{matrix}\right)\left(\begin{matrix}
 y'&z'\\
 0&0
\end{matrix}\right)=\left(\begin{matrix}
1&0
 \end{matrix}\right)\left(\begin{matrix}
 y&z\\
 0&0
\end{matrix}\right)\left(\begin{matrix}
 0&-1\\
 1&0
\end{matrix}\right)+\left(\begin{matrix}
1&0
 \end{matrix}\right)\left(\begin{matrix}
 0&-f(x)\\
 0&0
\end{matrix}\right)"

It is enough to solve the problem: "A'=A\sigma+F", where "A=\left(\begin{matrix}
 y&z\\
 0&0
\end{matrix}\right)", "\sigma=\left(\begin{matrix}
 0&-1\\
 1&0
\end{matrix}\right)" "F=\left(\begin{matrix}
 0&-f(x)\\
 0&0
\end{matrix}\right)". The solution of equation "A'=A\sigma" is "A=Ce^{\sigma t}" with matrix "C". We suppose that "C=C(t)". We substitute and get: "C'e^{\sigma t}+Ce^{\sigma t}{\sigma}=Ce^{\sigma t}{\sigma}+F". We get: "C=\int Fe^{-\sigma t}dt+K" with a constant matrix "K=\left(\begin{array}{cc}k_{11}&k_{12}\\k_{21}&k_{22}\end{array}\right)." Thus, "A=(\int Fe^{-\sigma t}dt+K)e^{\sigma t}". We notice that "y=\left(\begin{matrix}
1&0
 \end{matrix}\right)\left(\begin{matrix}
 y&z\\
 0&0
\end{matrix}\right)\left(\begin{matrix}
 1\\
 0
\end{matrix}\right)" . Thus, "y=\left(\begin{matrix}
1&0
 \end{matrix}\right)(\int Fe^{-\sigma t}dt+K)e^{\sigma t}\left(\begin{matrix}
 1\\
 0
\end{matrix}\right)". From the latter we can find the Green function.

Remind that "e^{\sigma t}=1+\sigma t+\frac{(\sigma t)^2}{2!}+\frac{(\sigma t)^3}{3!}+..." Compute powers of matrix "\sigma": "\sigma^2=\sigma=\left(\begin{matrix}
 -1&0\\
 0&-1
\end{matrix}\right)=-I". Thus, "e^{-\sigma t}=I(1-\frac{(-t)^2}{2!}+\frac{(-t)^4}{4!}).-+...+\sigma(-t-\frac{(-t)^3}{3!}+\frac{(-t)^5}{5!})+...=Icos(\,-t)+\sigma\,sin(-\,t)"

As a result, we get:"y=\left(\begin{matrix}
1&0
 \end{matrix}\right)(\int Fe^{-\sigma t}dt+K)e^{\sigma x}\left(\begin{matrix}
 1\\
 0
\end{matrix}\right)=\left(\begin{matrix}
1&0
 \end{matrix}\right)(\int \left(\begin{matrix}
 0&-f(t)\\
 0&0
\end{matrix}\right)\left(\begin{matrix}
 cos\,t&sin\,t\\
 -sin\,t&cos\,t
\end{matrix}\right)dt+\left(\begin{array}{cc}k_{11}&k_{12}\\k_{21}&k_{22}\end{array}\right))e^{\sigma x}\left(\begin{matrix}
 1\\
 0
\end{matrix}\right)=\left(\begin{matrix}
1&0
 \end{matrix}\right) \left(\begin{matrix}
 \int f(t)\,sin\,t \,dt+k_{11}&-\int f(t)\,cos\,t\,dt+k_{12}\\
 k_{21}&k_{22}\end{matrix}\right)\left(\begin{matrix}
 cos\,x&-sin\,x\\
 sin\,x&cos\,x
\end{matrix}\right)\left(\begin{matrix}
 1\\
 0
\end{matrix}\right)=\left(\begin{matrix}
1&0
 \end{matrix}\right)A\left(\begin{matrix}
 1\\
 0
\end{matrix}\right)=\left(\begin{matrix}
1&0
 \end{matrix}\right)\left(\begin{matrix}
 a_{11}&a_{12}\\
 a_{21}&a_{22}
\end{matrix}\right)\left(\begin{matrix}
 1\\
 0
\end{matrix}\right)=a_{11}".

"a_{11}= cos\,x(\int f(t)\,sin\,t \,dt+k_{11})+sin\,x(-\int f(t)\,cos\,t\,dt+k_{12})=k_{11} cos\,x+k_{12}sin\,x+\int f(t)(cos\,x\,sin\,t-sin\,xcos\,t) \,dt"

As we can see, "k_{11} cos\,x+k_{12}sin\,x" is a solution of the homogenous problem "y''+y=0". Therefore, we set "k_{11}=k_{12}=0" and receive that the Green function is: "G(x,t)=cos\,x\,sin\,t-sin\,xcos\,t"